Jim is k years old. Joe is k2 (k squared) years old. In 8 years Joe will be twice as old as Jim is then. How old is Joe now? NEeds to be done as a "Zero Product Property"
0 = k+8(2)

Answer by kapilsinghi(68)

Jim is k years old. Joe is k2 (k squared) years old. In 8 years Joe will be twice as old as Jim is then. How old is Joe now? NEeds to be done as a "Zero Product Property"
Jim Age after 8 years = k+8
Joe age after 8 years = k^2 +8
according to the condition "In 8 years Joe will be twice as old as Jim is then"
(k+8)*2= k^2 + 8
2k+16= k^2 + 8
0= k^2 -2k + 8 -16
0= k^2 -2k -8
0= k^2 -4k +2k -8
0= k(k-4)+2(k-4)
0= (k+2)(k-4)
k=-2 rejected
k=4
age of joe = k^2 = 4^2 = 16